OpenStudy (anonymous):

Determine the slope and the y-intercept of the graph of the equation. 6x - 4y = 24

OpenStudy (mertsj):

Solve for y y=3/2x-6

OpenStudy (anonymous):

3/2 slope -6 y-intercept

OpenStudy (mertsj):

Slope = 3/2 y=-6

OpenStudy (anonymous):

m will be your slope if you match the form: \[y = mx + b\] \[x - 4y = 24\] *bringng y to one side \[4y = 6x -24\] *divide out the 4 \[y = (6/4)x -6\] *simply faction \[y = (3/2)x -6\] m = 3/2